Q3 (5pt) A particle of mass m is attracted to a force center with the force of magnitude k/r2. Use the plane polar coordinates (r, ?) and the Hamiltonian method to find the equation of motion for r and ?.
Answer #1

The expression for the kinetic energy of the particle in polar coordinates is, T = 2m(32 +r??) The expression for the attract= = m(j? +r?0?) + The expression for the Hamiltonian of the particle is, H = ip, +ope- The expression for r polar coordinatei-P; m The expression for a polar coordinate of the particles momentum is, al PA = Po ao Then write the expression for deriv+ =ip. +0p{m{x+r20)) =()+(?).-10-17)2) 2mr2 PIE 2m 2mr r The equation of motion of the particle in position vector is, H = op2m = 2P: +0-0 i= (or) pe = mi The equation of motion of the particle in momentum vector is, . -a Para Or Calculate the equatiThe equation of motion of the particle in direction vector is, 0-OH op Calculate the equation of motion of the particle in diCalculate the equation of motion of the particle in momentum vector as follows: -OH ao 2mr2 Do = 0